Solve.

The length of a rectangle is 8 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the width of the rectangle.

Answer by edjones(8007)

Let L=length, w=width
L=2w+8
2L+2w=P
2(2w+8)+2w=34
4w+16+2w=34
6w+16-16=34-16
6w=18
6w/6=18/6
w=3
L=2*3+8
L=14
Ed